OpenStudy (anonymous):

3) A textbook company keeps track of its daily cost for printing whole textbooks to different schools but can only print up to 500 books a day. The daily cost (C) to print x number of textbooks is given by C(x) = 3.50x + 1200. Using the function, determine one possible daily cost by evaluating the function for any input of your choice.

OpenStudy (anonymous):

Interpret the statement C(20) = 1270. What does the 20 represent? What does the 1270 represent? What does the entire equation represent?

OpenStudy (anonymous):

Interpret the statement C(w) = p. What does the w represent? What does the p represent? What does the entire equation represent?

OpenStudy (anonymous):

Explain whether C(40.5) = 1341.75 is correct and possible given the context of this problem. Hint: Can you make .5 of a textbook?

OpenStudy (anonymous):

can't figure this one out.

OpenStudy (aum):

C(x) = 3.50x + 1200 ---- (1) where C is the cost and x is the number of textbooks printed per day. The maximum value x can have is 500. "Using the function, determine one possible daily cost by evaluating the function for any input of your choice." Let us find the cost to print 100 textbooks per day. x = 100. Put that in (1): C(100) = 3.5 * 100 + 1200 = 350 + 1200 = $1550. Therefore it will cost $1,550 to print 100 textbooks per day.

OpenStudy (aum):

Interpret the statement C(20) = 1270. What does the 20 represent? What does the 1270 represent? What does the entire equation represent? What does the 20 represent? x = 20. Therefore 20 represents the number of textbooks printed per day. What does the 1270 represent? $1270 represents the cost to print 20 textbooks per day. The entire equation, C(20) = 1270, tells us the cost to print 20 textbooks per day is $1270.

OpenStudy (anonymous):

thank you for the help, I was on the problem for more than 30 minutes